OpenStudy (anonymous):

Sam is observing the velocity of a car at different times. After two hours, the velocity of the car is 50 km/h. After six hours, the velocity of the car is 54 km/h. Part A: Write an equation in two variables in the standard form that can be used to describe the velocity of the car at different times. Show your work and define the variables used. (5 points) Part B: How can you graph the equation obtained in Part A for the first seven hours? (5 points)

OpenStudy (anonymous):

I need help on Part A...

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

You will need some idea about velocity. This one seems to be changing over time, so what functions can you think of that increase over time?

OpenStudy (anonymous):

I have from looking at similar questions that v=2t+50 but idk how they got it

OpenStudy (anonymous):

I know it will be linear if that has to do with anything but idk how to get a function from the question.

OpenStudy (tkhunny):

Super. Generalize it and you'll be almost done. Let's go with the linear function. v(t) = at + b How's that? Can you find 'a' and 'b' such that v(2) = 50 and v(6) = 54?

OpenStudy (anonymous):

Im still confused as to how to find what each variable is in each thing.

OpenStudy (anonymous):

@pooja195 can you help??

OpenStudy (tkhunny):

You must examine the equation. I'll do one. We have v(t) = at + b We know v(2) = 50 This means v(2) = a(2) + b = 50 Now, what do we know about v(6)?

OpenStudy (anonymous):

That v(6)= a(6)+b =54

OpenStudy (anonymous):

@pooja195 help